# LLMs as Writing Assistants: Exploring Perspectives on Sense of Ownership and Reasoning

### v2

@@ Abstract.
Sense of ownership in writing confines our investment of thoughts, time, and contribution, leading to attachment to the output. However, using writing assistants introduces a mental dilemma, as some content isn’t directly our creation. For instance, we tend to credit Large Language Models (LLMs) more in creative tasks, even though all tasks are equal for them. Additionally, while we may not claim complete ownership of LLM-generated content, we freely claim authorship. We conduct a short survey to examine these issues and understand underlying cognitive processes in order to gain a better knowledge of human-computer interaction in writing and improve writing aid systems.
@@ 1. Introduction
Sense of Ownership in Writing. Ownership in writing transcends the physical act of possessing pen and paper; it’s a profound bond between thoughts and words. Crafting sentences and paragraphs involves investing in ourselves, nurturing a sense of ownership over our ideas. It’s more than claiming credit; it’s cultivating a relationship with our creation, akin to tending a cherished plant. Each pen stroke embodies our identity, manifesting thoughts and emotions onto the page. This ownership empowers us to share our stories, knowing they carry our essence. In content writing, this sense of ownership arises from a complex interplay of cognitive and emotional experiences (Eli, 2013), weaving personal commitment and connection into every word, whether it’s an assignment, a birthday wish, or an article.
Sense of Ownership in LLM Assisted Writing. LLMs are becoming increasingly popular in both education and business because of their exceptional performance in a wide range of applications (Guo et al., 2023). LLMs can be used effectively in different writing tasks like creative stories and essays (Huang et al., 2023; Yuan et al., 2022; Chung et al., 2022), academic writing (Bekker, 2024; Jarrah et al., 2023), and legal write-ups (Nay et al., 2023). When utilizing LLMs like ChatGPT, Gemini or Microsoft Copilot, the sense of ownership in writing may evolve, as the process becomes collaborative, blending human input with machine-generated content. Additionally, LLMs can facilitate the exploration of diverse perspectives and styles, altering the traditional notions of authorship and creativity in content creation.
LLMs serve both creative and non-creative tasks, prompting the question: do our feelings of ownership align similarly across both creative and non-creative tasks when working with LLMs? As humans, we often feel more attached to creative tasks, attributing a greater sense of ownership to them (Cassidy et al., 2013). Another focus is investigating the link between a sense of ownership and authorship in LLM-based writing assistants, which is crucial, especially given the prevalent issue of plagiarism among humans (Park, 2017; Milano et al., 2023). Also, different social dynamics are involved here too, as indicated by Gero et al. (Gero et al., 2023). Understanding how individuals perceive ownership and authorship in the context of LLM assistance can help address concerns about originality and attribution in content creation.
@@
Understanding the nuances in our perception of ownership and authorship highlights the importance of addressing them when developing writing assistants. We conducted a brief survey with 35 participants to gain insight into and explore their perspectives to try to understand these complexities better. By understanding these, we hope to create more effective tools that align with users’ needs and enhance their writing experiences. Additionally, navigating these complexities can help us use LLMs more efficiently in real-life scenarios, ensuring they serve as valuable aids rather than obstacles in the creative process.
@@ 2. Exploring Problems and Perspectives
We conducted a brief survey with 35 participants to evaluate their sense of ownership while using writing assistants, and the questions that arose above. Survey questions are mentioned in Section A. The demographic profile of the participants are available in Section B.
2.1. Type of Content and Ownership
2.2. Authorship and Ownership
Draxler et al. (Draxler et al., 2024) also acknowledged and investigated this phenomenon and named it AI Ghostwriter Effect, potentially influencing behaviors related to ghostwriting111 Ghostwriting is the practice of writing content on behalf of someone else who takes credit for the work, often without any acknowledgment of the ghostwriter’s involvement. (Rhodes, 2000), a common thing in writing industry. This shows that there is a gap between acknowledging ownership and attributing authorship, which may have its roots in a desire for self-recognition or in the belief that authorship involves more than just possession—such as presentation or accountability. Though this is connected to our second research question, but our work distinctly focuses on disentangling the perceptions of ownership and authorship when using LLM writing assistants, a nuanced aspect not directly explored in previous work on the AI Ghostwriter Effect. Our methodology is designed to explicitly probe this ownership-authorship dichotomy through carefully constructed experiments and surveys. This targeted approach allows us to deepen our understanding of human-computer writing interactions.
Figure 2. Survey Statistics: Different Levels of Sense of Ownership
After reminding participants of their role in providing the prompts (as discussed in Section A), a significant shift occurred: more participants claimed ownership of the content (’Yes’: increased from 17.1% to 28.6%; ’Maybe’: increased from 31.4% to 48.6%; ’No’: reduced from 51.4% to 22.9%, as shown in Figure 2 (c) ). So, reminding participants of their role in providing prompts likely increased their sense of ownership by highlighting their contribution to the content creation process, leading to a reevaluation of their ownership perception. After reminding them about editing the LLM response, the sense of authorship surged, with 62.9% of participants claiming ownership of the content (Figure 2 (d)).
3. Discussion
We acknowledge a gap in content types, as only two types (general and creative) are explored. Future work in these areas should extend the content types.
In this work, we aim to establish the foundation for the creation of increasingly complicated and effective writing assistance systems, as well as to clarify the complexity of human-computer interaction in writing activities. To achieve this, it is essential to enhance transparency and accountability in LLM-generated content. Providing clear guidelines and standards for disclosing the involvement of LLMs in content creation can help individuals make informed judgments about ownership. Collaborative approaches that emphasize human-AI partnerships, rather than relying solely on AI-generated content, may also help reinforce a sense of ownership and authorship among individuals involved in creative tasks.
4. Ethics and Laws
Sense of ownership is intricately linked to tangible ownership, ethical considerations, and consequently intersects with legal frameworks and jurisdictions. And, this issue regarding AI-generated content remains complex and varies across jurisdictions. In the US, current law typically requires human authorship for ownership, as seen in the Thaler v. Perlmutter case. However, a Chinese court has ruled differently, granting AI-generated content protection under copyright law. Legal decisions and ongoing discussions shape our understanding of authorship and ownership in the AI era. Courts would likely examine how LLMs learn and produce content, potentially concluding that AI, not humans, authored the output, thus impacting ownership rights (Draxler et al., 2024; Tombekai, 2020). For instance, OpenAI retains ownership of the language model and the content generated by ChatGPT, while users are acknowledged as the collaborators but not the owners or authors of the AI-generated text (Draxler et al., 2024). This legal ambiguity may influence individuals’ actual of ownership over AI-generated content, as they grapple with the blurred lines between human input and machine output, potentially challenging traditional notions of ownership in the digital age.
5. Conclusion
This study explores into two prevalent mental dilemmas surrounding LLM-assisted writing: firstly, the discrepancy in crediting LLMs across various types of content, influencing individuals’ sense of ownership; and secondly, the interplay between authorship claims and the perceived ownership of LLM-generated content. Through survey data analysis, we illustrate the prevalence of these dilemmas and the underlying thought processes. By addressing these complexities, our research aims to pave the path toward a deeper understanding of human-computer interaction dynamics in writing tasks and foster the development of more effective writing assistance systems.
Below is a brief overview of conclusions supported by data regarding participants’ motives, alongside some speculative insights:
  • Our research shows that the sense of ownership differs with the type of content when utilizing LLMs as writing assistants. In creative endeavors like poems and stories, individuals are less likely to claim ownership and are inclined to credit LLMs. Conversely, in general content creation, individuals feel a stronger sense of ownership over the produced content.
  • Our research highlights an interesting interplay between ownership and authorship perceptions when utilizing LLMs as writing assistants, revealing that individuals may not assert ownership of LLM-generated content but demonstrate a willingness to claim authorship.
  • The experiments revealed a progression in ownership perception: participants were initially hesitant but willing to attribute authorship. Reminders of their role in providing prompts boosted ownership perception, and further engagement through editing the LLM response notably increased the sense of authorship, showcasing the impact of active involvement on ownership perception.
Appendix A Survey Questions
@@
Appendix B Demographic Profile of Survey Participants

### v4

@@ Abstract.
Sense of ownership in writing confines our investment of thoughts, time, and contribution, leading to attachment to the output. However, using writing assistants introduces a mental dilemma, as some content isn’t directly our creation. For instance, we tend to credit Large Language Models (LLMs) more in creative tasks, even though all tasks are equal for them. Additionally, while we may not claim complete ownership of LLM-generated content, we freely claim authorship. We conduct a short survey to examine these issues and understand underlying cognitive processes in order to gain a better knowledge of human-centered aspects in writing and improve writing aid systems.
@@ 1. Introduction
Figure 1. Mental dilemmas while using LLMs as Writing Assistants: Who wrote it? Who owns it?
LLMs serve both creative and non-creative tasks, prompting the question: do our feelings of ownership align similarly across both creative and non-creative tasks when working with LLMs? As humans, we often feel more attached to creative tasks, attributing a greater sense of ownership to them (Cassidy et al., 2013). Another focus is investigating the link between a sense of ownership and authorship in LLM-based writing assistants, which is crucial, especially given the prevalent issue of plagiarism among humans (Park, 2017; Milano et al., 2023). Also, different social dynamics are involved here too, as indicated by Gero et al. (2023). Understanding how individuals perceive ownership and authorship in the context of LLM assistance can help address concerns about originality and attribution in content creation.
@@
Understanding the complexities in our perception of ownership and authorship highlights the importance of addressing them when developing writing assistants. We conducted a brief survey with 35 participants to gain insight into and explore their perspectives to try to understand these complexities better. By understanding these, we hope to create more effective tools that align with users’ needs and enhance their writing experiences. Additionally, navigating these complexities can help us use LLMs more efficiently in real-life scenarios, ensuring they serve as valuable aids rather than obstacles in the creative process.
2. Cognition, Ownership, and the Rise of Language Models
2.1. Sense of Ownership in Writing.
Ownership in writing transcends the physical act of possessing pen and paper; it’s a profound bond between thoughts and words. Crafting sentences and paragraphs involves investing in ourselves, nurturing a sense of ownership over our ideas. It’s more than claiming credit; it’s cultivating a relationship with our creation, akin to tending a cherished plant. Each pen stroke embodies our identity, manifesting thoughts and emotions onto the page. This ownership empowers us to share our stories, knowing they carry our essence. In content writing, this sense of ownership arises from a complex interplay of cognitive and emotional experiences (Eli, 2013), weaving personal commitment and connection into every word, whether it’s an assignment, a birthday wish, or an article.
$2.$ Sense of Ownership in LLM Assisted Writing.
LLMs are becoming increasingly popular in both education and business because of their exceptional performance in a wide range of applications (Guo et al., 2023). LLMs can be used effectively in different writing tasks like creative stories and essays (Huang et al., 2023; Yuan et al., 2022; Chung et al., 2022), academic writing (Bekker, 2024; Jarrah et al., 2023), and legal write-ups (Nay et al., 2023). When utilizing LLMs like ChatGPT, Gemini or Microsoft Copilot, the sense of ownership in writing may evolve, as the process becomes collaborative, blending human input with machine-generated content. Additionally, LLMs can facilitate the exploration of diverse perspectives and styles, altering the traditional notions of authorship and creativity in content creation.
@@ 3. Exploring Problems and Perspectives
We conducted a brief survey with 35 participants to evaluate their sense of ownership while using writing assistants, and the questions that arose above. The demographic profile of the participants are available in Appendix A.
3.1. Survey Questions
@@
3.2. Type of Content and Ownership
$3.$ Authorship and Ownership
Draxler et al. (2024) also acknowledged and investigated this phenomenon and named it AI Ghostwriter Effect, potentially influencing behaviors related to ghostwriting111 Ghostwriting is the practice of writing content on behalf of someone else who takes credit for the work, often without any acknowledgment of the ghostwriter’s involvement., a common thing in writing industry (Rhodes, 2000). This shows that there is a gap between acknowledging ownership and attributing authorship, which may have its roots in a desire for self-recognition or in the belief that authorship involves more than just possession—such as presentation or accountability. Though this is connected to our second research question, but our work distinctly focuses on disentangling the perceptions of ownership and authorship when using LLM writing assistants, an interesting aspect not directly explored in previous work. Our methodology is designed to explicitly probe this ownership-authorship dichotomy through carefully constructed experiments and surveys. This targeted approach allows us to deepen our understanding of human-centric aspects of LLMs.
Figure 3. Survey Statistics: Different Levels of Sense of Ownership
After reminding participants of their role in providing the prompts (as discussed in Section 3.1), a significant shift occurred: more participants claimed ownership of the content (’Yes’: increased from 17.1% to 28.6%; ’Maybe’: increased from 31.4% to 48.6%; ’No’: reduced from 51.4% to 22.9%, as shown in Figure 3 (c) ). So, reminding participants of their role in providing prompts likely increased their sense of ownership by highlighting their contribution to the content creation process, leading to a reevaluation of their ownership perception. After reminding them about editing the LLM response, the sense of authorship surged, with 62.9% of participants claiming ownership of the content (Figure 3 (d)).
4. Discussion
Our goal is to lay the groundwork for better writing assistance systems and improve our understanding of human-centered aspects of writing. To achieve this, we need to increase transparency and accountability in content created with LLMs. Clear guidelines for disclosing LLM involvement can help people make informed decisions about ownership. Emphasizing collaborative approaches that blend human and AI contributions in creative tasks can strengthen feelings of ownership and authorship among individuals involved.
Limitations. One limitation of this study is the small sample size, consisting of only 35 participants. Increasing the number of participants could lead to more robust and meaningful research outcomes. We also acknowledge a gap in content types, as only two types (general and creative) are explored. Future work in these areas should extend the content types and LLM application areas.
5. Ethics and Laws
Sense of ownership is intricately linked to tangible ownership, ethical considerations, and consequently intersects with legal frameworks and jurisdictions. And, this issue regarding AI-generated content remains complex and varies across jurisdictions. In the US, current law typically requires human authorship for ownership, as seen in the Thaler v. Perlmutter case. However, a Chinese court has ruled differently, granting AI-generated content protection under copyright law. Legal decisions and ongoing discussions shape our understanding of authorship and ownership in the AI era. Courts would likely examine how LLMs learn and produce content, potentially concluding that AI, not humans, authored the output, thus impacting ownership rights (Draxler et al., 2024; Tombekai, 2020). For instance, OpenAI retains ownership of the language model and the content generated by ChatGPT, while users are acknowledged as collaborators but not the owners or authors of the AI-generated text (Draxler et al., 2024). This legal ambiguity may influence individuals’ actual sense of ownership over AI-generated content as they grapple with the blurred lines between human input and machine output, potentially challenging traditional notions of ownership in the digital age.
6. Findings and Conclusion
This study explores into two prevalent mental dilemmas surrounding LLM-assisted writing: firstly, the discrepancy in crediting LLMs across various types of content, influencing individuals’ sense of ownership; and secondly, the interplay between authorship claims and the perceived ownership of LLM-generated content. Below is a brief overview of conclusions supported by data regarding participants’ motives, alongside some speculative insights:
  • Our research indicates that ownership perceptions vary based on content type when using LLMs as writing assistants. In creative works such as poems and stories, individuals tend to credit LLMs and feel less ownership. Conversely, in general content creation, individuals exhibit a stronger sense of ownership over the produced content.
  • Our research reveals a interesting relationship between ownership and authorship perceptions when using LLMs as writing assistants. Individuals may not assert ownership over LLM-generated content but show a willingness to claim authorship.
  • The experiments demonstrated a progression in ownership perception: participants were initially hesitant but willing to claim authorship. Reminders of their role in providing prompts increased ownership perception, and further engagement through editing the LLM response notably enhanced authorship perception, illustrating the impact of active involvement on ownership.
Through survey data analysis, in this work, we aim to identify prevalent dilemmas and underlying thought processes in human-centric aspects of LLM usage to advance understanding and enhance writing assistance system development.
Author Contributions
ATW conceived the core idea, developed the methodology, designed the experiments and the data collection, performed some experiments, conducted all the analyses, prepared visualizations, wrote and edited core parts of the paper, and led the whole project. MRI collected the data, formal analysis and was involved in initial draft writing and editing. RI helped in background study, writing and editing.
Appendix A Demographic Profile of Survey Participants